OpenStudy (anonymous):

solve by elimination method 7r-5s=16 5r+7s=96

OpenStudy (anonymous):

Multiply top equation by 5 and bottom equation by 7: 35r-25s=80 35r+49s=672 Now subtract the second equation from the first: -74s=-592 Solving: s=8 Now input s into either of the original equations and solve for r. I get r=8.

OpenStudy (anonymous):

7r-5s=16 multiply by 7 5r+7s=96 multiply by 5

OpenStudy (anonymous):

so the answer would be 8?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is their more then one solution too?

OpenStudy (anonymous):

only one solution. r=s=8

OpenStudy (anonymous):

so what is the ordered pair?

OpenStudy (anonymous):

fractions or integers?

OpenStudy (anonymous):

Uh...these are integers. If you want to write the solution as an ordered pair then (r,s)=(8,8)

OpenStudy (anonymous):

There is no natural "order" unless you provide one. By that I mean it doesn't matter whether u write r first or s first, just so long as u make it clear which is which (in this cvase, it doesn't matter since they are both 8).